OpenStudy (anonymous):

find a formula for the inverse of the function ______ f(x) = √10-3x it would be really helpful to explain how you got the answer :)

OpenStudy (dumbcow):

y = sqrt(10-3x) now try to solve for x square both sides y^2 = 10-3x y^2 -10 = -3x (y^2 -10)/-3 = x or if you distribute the neg x = (10-y^2)/3

OpenStudy (anonymous):

tyvm :)